# On characterization and existence of a constrained correlated equilibria in Markov games

### v1

@@ 1 Introduction
Let $\mathbb{N}$ and $\mathbb{R}$ denote the sets of natural and real numbers, respectively. For any $m\in\mathbb{N}$, we define $[m]:=\{1,\dots,m\}$. We denote by $x_{m:n}$ the sequence $\{x_{m},x_{m+1},\dots,x_{n}\}$. Let $\mathbf{1}_{n}$ denote the all-ones vector of dimension $n$. Given a finite set $\mathcal{X}$, we denote the indicator function by $\mathds{1}_{\mathcal{X}}(x)$. Furthermore, the probability simplex over $\mathcal{X}$ is denoted by $\Delta({\mathcal{X}})$ and the cardinality of $\mathcal{X}$ is denoted by $|\mathcal{X}|$. Given any two finite sets $\mathcal{X}$ and $\mathcal{Y}$ in $2^{\mathbb{R}^{d}}$, we define the Minkowski sum of the sets $\mathcal{X}$ and $\mathcal{Y}$ as $\mathcal{X}+\mathcal{Y}=\{x+y\mid x\in\mathcal{X},\,y\in\mathcal{Y}\}$. Given a point-to-set mapping $f:\mathcal{X}\to 2^{\mathcal{Y}}$, we say that $f$ is upper semi-compact if it is upper semi-continuous and $f(x)$ is compact for all $x\in\mathcal{X}$ [5, Definition 2.2].
@@ 2 A finite-horizon Markov game with coupling constraints
[…]
@@ 4 Existence of a constrained correlated equilibrium
with constraint thresholds $c^{1,1}=\frac{1}{2}$ and $c^{2,1}=\frac{1}{3}$. It is straightforward to verify that the strong Slater’s condition (Assumption 4.1) does not hold, as there exists no feasible modification for either $P1$ or $P2$ when the policy is $\pi((a^{1},a^{2}))=\mathds{1}_{(a^{1},a^{2})=(2,2)}$. Furthermore, no constrained correlated equilibrium exists in this game since for any feasible policy $\pi$, if $P2$ applies its best $i$-feasible modification with $i=2$, the resulting modified policy becomes $i$-infeasible with $i=1$ for $P1$, thereby violating the equilibrium condition. A detailed mathematical justification is provided in Appendix B.1.
@@
For any player $i\in\mathcal{N}$ and any policy $\pi\in\Pi_{M}$ on the boundary of $\mathcal{C}_{\pi}$, i.e., $\pi\in\{\pi\in\Pi_{M}\mid\exists j\in[J]\,\,s.t.\,\,V^{g^{j}}(\pi)=c^{j}\}$, at least one of the following conditions hold:
  1. 1.
    There exists a modification $\phi^{i}\in\Phi^{i}_{M}$ such that $V^{g^{j}}(\phi^{i}\circ\pi)>c^{j},\,\forall j\in[J].$
  2. 2.
    1. (a)
      For all $j\in[J]$, it holds that: $\min_{\phi^{i}\in\Phi^{i}_{M}}V^{g^{j}}(\phi^{i}\circ\pi)<c^{j}$.
    2. (b)
      There exists an $\alpha\in\Delta(K^{i})$ with positive weights, namely $\alpha_{k}>0$ for all $k\in[K^{i}]$, such that:
@@
Note that the above assumption indeed weakens Assumption 4.1. Specifically, the first condition only requires existence of a strictly feasible modification for policies on the boundary of the feasible set $\mathcal{C}_{\pi}$, rather than for all policies $\Pi_{M}$ as required by the strong Slater’s condition. The second condition complements the first and imposes an additional requirement in (b). For illustration, in normal-form games, (b) is equivalent to the existence of a feasible modification in the interior of $\Delta(\mathcal{A}^{i})$. The next example shows that there exists a Markov game with common coupling constraints admitting a constrained correlated equilibrium, in which the second condition of the above assumption is satisfied, but strong Slater’s condition is violated.
@@
with constraint thresholds $c^{j}=\frac{1}{4}$ for all $j\in[4]$. In this setting, the feasible policy set $\mathcal{C}_{\pi}$ contains only one policy, namely the uniform distribution $\pi(a)=\frac{1}{4}$ for all $a\in\mathcal{A}$. As the interior of $\mathcal{C}_{\pi}$ is empty, there exists no modification $\phi^{i}$ for all $i\in[2]$, such that $\phi^{i}\circ\pi$ lies in the interior of $\mathcal{C}_{\pi}$. Consequently, the first condition of Assumption 4.2 fails to hold. This, in turn, also implies that the strong Slater’s condition is violated. However, the second condition of Assumption 4.2 is satisfied and it is straightforward to verify that the unique feasible policy $\pi$ is indeed a constrained correlated equilibrium. We provide a detailed mathematical justification in Appendix B.2.
@@ 5.1 Preliminary results
$\max_{\alpha}$ $\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi)$
@@ A.1.2 Properties of the first MDP
×$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})\phi^{i}_{t}(\hat{a}_{t}^{i}|(s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i}))$
=$\overset{(i)}{=}$ $\sum_{s^{\prime}_{1:t-1},a^{\prime}_{1:t-1}}\sum_{s_{1:t-2},a_{1:t-2}}\sum_{s_{t-1},a_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
×$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
=$\overset{(ii)}{=}$ $\frac{1}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(\hat{a}_{t-1}^{i},a_{t-1}^{-i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
×$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i}).$ (15)
@@
×$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
=$\overset{(i)}{=}$ $\frac{\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(\hat{a}_{t-1}^{i},a_{t-1}^{-i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
×$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\bar{\phi}^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
=$\overset{(ii)}{=}$ $\frac{\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(a_{t-1}^{-i},\hat{a}_{t-1}^{i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
×$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
=$\overset{(iii)}{=}$ ${\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}\left(\sum_{\hat{a}_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\right).$ (16)
@@
=$\overset{(i)}{=}$ $|\mathcal{A}^{i}|\sum_{a_{1}^{i}}\tilde{\rho}(s_{1},a_{1}^{i})\phi^{i}_{1}(\hat{a}^{i}_{1}|s_{1},a_{1}^{i})\pi_{1}((a_{1}^{i},a_{1}^{-i})\mid s_{1})$
=$\overset{(ii)}{=}$ $\sum_{a_{1}^{i}}\rho(s_{1})\phi_{1}^{i}(\hat{a}^{i}_{1}|s_{1},a_{1}^{i})\pi_{1}((a_{1}^{i},a_{1}^{-i})\mid s_{1})$
@@
=$\overset{(i)}{=}$ $\sum_{\bar{s}_{1:t-2},\bar{a}_{1:t-2}}\sum_{\bar{s}_{t-1}}\sum_{\bar{a}_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})$
×$\times\tilde{P}_{t-1}^{\pi}(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}|(\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})$
=$\overset{(ii)}{=}$ $|\mathcal{A}^{i}|^{-1}\sum_{\bar{s}_{1:t-2},\bar{a}_{1:t-2}}\sum_{\bar{s}_{t-1}}\sum_{\bar{a}_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})\pi_{t-1}((\bar{a}_{t-1}^{i},\bar{a}_{t-1}^{-i})\mid\bar{s}_{t-1})$
×$\times P_{t-1}(s_{t}|\bar{s}_{t-1},(\hat{a}_{t-1}^{i},\bar{a}_{t-1}^{-i}))\mathds{1}_{a_{t-1}=(\hat{a}_{t-1}^{i},\bar{a}_{t-1}^{-i}),(s_{1:t-1},a_{1:t-2})=(\bar{s}_{1:t-1},\bar{a}_{1:t-2})}\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})$
=$\overset{(iii)}{=}$ $|\mathcal{A}^{i}|^{-1}\sum_{\bar{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},\bar{a}_{t-1}^{i}),a_{t-1}^{i})\pi_{t-1}((\bar{a}_{t-1}^{i},a_{t-1}^{-i})\mid s_{t-1})P_{t-1}(s_{t}|s_{t-1},a_{t-1})$
×$\times\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}).$
@@
×$\times P_{t-1}(s_{t}|s_{t-1},a_{t-1})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
=$\overset{(i)}{=}$ $\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}d_{t-1}^{\phi^{i}\circ\pi}(s_{1:t-1},a_{1:t-1})P_{t-1}(s_{t}|s_{t-1},a_{t-1})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$

### v2

# On characterization and existence of constrained correlated equilibria in Markov games

@@ 1 Introduction
Let $\mathbb{N}$ and $\mathbb{R}$ denote the sets of natural and real numbers, respectively. For any $m\in\mathbb{N}$, we define $m]:=\{1,\dots,m\}$. We denote by $x_{m:n}$ the sequence $\{x_{m},x_{m+1},\dots,x_{n}\}$. Let $\mathbf{1}_{n}$ denote the all-ones vector of dimension $n$. Given a finite set $\mathcal{X}$, we denote the indicator function by $\mathds{1}_{\mathcal{X}}(x)$. Furthermore, the probability simplex over $\mathcal{X}$ is denoted by $\Delta({\mathcal{X}})$ and the cardinality of $\mathcal{X}$ is denoted by $|\mathcal{X}|$. Given any two finite sets $\mathcal{X}$ and $\mathcal{Y}$ in $2^{\mathbb{R}^{d}}$, we define the Minkowski sum of the sets $\mathcal{X}$ and $\mathcal{Y}$ as $\mathcal{X}+\mathcal{Y}=\{x+y\mid x\in\mathcal{X},\,y\in\mathcal{Y}\}$. Given a point-to-set mapping $f:\mathcal{X}\to 2^{\mathcal{Y}}$, we say that $f$ is upper semi-compact if it is upper semi-continuous and $f(x)$ is compact for all $x\in\mathcal{X}$ [5, Definition 2.2].
@@ 2 A finite-horizon Markov game with coupling constraints
where $\mathcal{N}=[N]$ denotes the set of players and $H\in\mathbb{N}$ denotes a finite horizon. The set $\mathcal{S}$ denotes a finite state space and each player $i\in\mathcal{N}$ has a finite action space $\mathcal{A}^{i}$. Let $a=(a^{1},\ldots,a^{N})\in\mathcal{A}$ denote the joint action profile, where $\mathcal{A}=\Pi_{i=1}^{N}A^{i}$ is the joint action space. Similarly, let $a^{-i}=(a^{1},\ldots,a^{i-1},a^{i+1},\ldots,a^{N})\in\mathcal{A}^{-i}=\Pi_{j\neq i}\mathcal{A}^{j}$ denote the joint action profile of all players except player $i$. Furthermore, each player $i\in\mathcal{N}$ has a reward function $r^{i}:=\{r^{i}_{t}\}_{t\in[H]}$, where $r^{i}_{t}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$. Additionally, each player has $J$ constraint functions $g^{i,j}:=\{g^{i,j}_{t}\}_{t\in[H]}$ for $j\in[J]$, where $g^{i,j}_{t}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$. The threshold with respect to each player $i$’s constraints $g^{i,j}$ is denoted by $c^{i,j}\in\mathbb{R}$. We refer to the set of constraints $\{g^{i,j}\}_{j\in[J]}$ as playerwise coupling constraints since each player has its own set of constraints that depend on the joint action profile. If, however, the constraint functions and threshold values are equal across all players, i.e., $g^{i,j}=g^{j}$ and $c^{i,j}=c^{j}$ for all $j\in[J]$ and $i\in\mathcal{N}$, then we refer to these types of constraints as common coupling constraints, as each player is subject to the same set of constraints. The transition kernel is denoted by $P:=\{P_{t}\}_{t=1\in[H-1]}$, where $P_{t}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$. Specifically, $P_{t}(s_{t+1}|s_{t},a_{t})$ denotes the probability of transitioning from state $s_{t}\in\mathcal{S}$ to state $s_{t+1}\in\mathcal{S}$ under the joint action $a_{t}\in\mathcal{A}$. Lastly, $\rho\in\Delta(\mathcal{S})$ denotes the initial state distribution.
@@ 4 Existence of a constrained correlated equilibrium
with constraint thresholds $c^{1,1}=\frac{1}{2}$ and $c^{2,1}=\frac{1}{3}$. It is straightforward to verify that the strong Slater’s condition (Assumption 4.1) does not hold, as there exists no feasible modification for either $P1$ or $P2$ when the policy is $\pi((a^{1},a^{2}))=\mathds{1}_{(a^{1},a^{2})=(2,2)}$. Furthermore, no constrained correlated equilibrium exists in this game since for any feasible policy $\pi$, if $P2$ applies its best $i$-feasible modification with $i=2$, the resulting modified policy strictly increases player $2$’s utility, thereby violating the equilibrium condition. A detailed mathematical justification is provided in Appendix B.1.
@@
For any player $i\in\mathcal{N}$ and any policy $\pi\in\Pi_{M}$ on the boundary of $\mathcal{C}_{\pi}$, i.e., $\pi\in\{\pi\in\Pi_{M}\mid\exists j\in[J]\,\,s.t.\,\,V^{g^{j}}(\pi)=c^{j}\}$, there exists an $\alpha\in\Delta(K^{i})$ with positive weights, namely $\alpha_{k}>0$ for all $k\in[K^{i}]$, such that:
@@
Note that the above assumption is indeed weaker than Assumption 4.1. Namely, it can easily be verified that Assumption 4.1 implies Assumption 4.2. The converse, however, is not true as the following example demonstrates. It shows that there exists a Markov game with common coupling constraints that admits a constrained correlated equilibrium, where the above assumption holds but the strong Slater’s condition is violated.
@@
with constraint thresholds $c^{j}=\frac{1}{4}$ for all $j\in[4]$. In this setting, the feasible policy set $\mathcal{C}_{\pi}$ contains only one policy, namely the uniform distribution $\pi(a)=\frac{1}{4}$ for all $a\in\mathcal{A}$. As the interior of $\mathcal{C}_{\pi}$ is empty, there exists no modification $\phi^{i}$ for all $i\in[2]$, such that $\phi^{i}\circ\pi$ lies in the interior of $\mathcal{C}_{\pi}$. Consequently, the strong Slater’s condition is violated. However, Assumption 4.2 is satisfied and it is straightforward to verify that the unique feasible policy $\pi$ is indeed a constrained correlated equilibrium. We provide a detailed mathematical justification in Appendix B.2.
@@ 5.1 Preliminary results
$\max_{\alpha}\$$\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi)$
@@ A.1.2 Properties of the first MDP
$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})\phi^{i}_{t}(\hat{a}_{t}^{i}|(s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i}))$
$\overset{(i)}{=}$ $\sum_{s^{\prime}_{1:t-1},a^{\prime}_{1:t-1}}\sum_{s_{1:t-2},a_{1:t-2}}\sum_{s_{t-1},a_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\overset{(ii)}{=}$ $\frac{1}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(\hat{a}_{t-1}^{i},a_{t-1}^{-i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i}).$ (15)
@@
$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\overset{(i)}{=}$ $\frac{\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(\hat{a}_{t-1}^{i},a_{t-1}^{-i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\bar{\phi}^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\overset{(ii)}{=}$ $\frac{\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(a_{t-1}^{-i},\hat{a}_{t-1}^{i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\overset{(iii)}{=}$ ${\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}\left(\sum_{\hat{a}_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\right).$ (16)
@@
$\overset{(i)}{=}$ $|\mathcal{A}^{i}|\sum_{a_{1}^{i}}\tilde{\rho}(s_{1},a_{1}^{i})\phi^{i}_{1}(\hat{a}^{i}_{1}|s_{1},a_{1}^{i})\pi_{1}((a_{1}^{i},a_{1}^{-i})\mid s_{1})$
$\overset{(ii)}{=}$ $\sum_{a_{1}^{i}}\rho(s_{1})\phi_{1}^{i}(\hat{a}^{i}_{1}|s_{1},a_{1}^{i})\pi_{1}((a_{1}^{i},a_{1}^{-i})\mid s_{1})$
@@
$\overset{(i)}{=}$ $\sum_{\bar{s}_{1:t-2},\bar{a}_{1:t-2}}\sum_{\bar{s}_{t-1}}\sum_{\bar{a}_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\times\tilde{P}_{t-1}^{\pi}(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}|(\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})$
$\overset{(ii)}{=}$ $|\mathcal{A}^{i}|^{-1}\sum_{\bar{s}_{1:t-2},\bar{a}_{1:t-2}}\sum_{\bar{s}_{t-1}}\sum_{\bar{a}_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})\pi_{t-1}((\bar{a}_{t-1}^{i},\bar{a}_{t-1}^{-i})\mid\bar{s}_{t-1})$
$\times P_{t-1}(s_{t}|\bar{s}_{t-1},(\hat{a}_{t-1}^{i},\bar{a}_{t-1}^{-i}))\mathds{1}_{a_{t-1}=(\hat{a}_{t-1}^{i},\bar{a}_{t-1}^{-i}),(s_{1:t-1},a_{1:t-2})=(\bar{s}_{1:t-1},\bar{a}_{1:t-2})}\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})$
$\overset{(iii)}{=}$ $|\mathcal{A}^{i}|^{-1}\sum_{\bar{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},\bar{a}_{t-1}^{i}),a_{t-1}^{i})\pi_{t-1}((\bar{a}_{t-1}^{i},a_{t-1}^{-i})\mid s_{t-1})P_{t-1}(s_{t}|s_{t-1},a_{t-1})$
$\times\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}).$
@@
$\times P_{t-1}(s_{t}|s_{t-1},a_{t-1})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$\overset{(i)}{=}$ $\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}d_{t-1}^{\phi^{i}\circ\pi}(s_{1:t-1},a_{1:t-1})P_{t-1}(s_{t}|s_{t-1},a_{t-1})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$